OpenStudy (lgbasallote):

Find the particular solution \[\large xy^2 dx + e^x dy = 0\] when \(x \rightarrow \infty\) and \(y \rightarrow \frac 12\) what do those approach mean -_-

OpenStudy (anonymous):

too hard, I quit.

OpenStudy (apoorvk):

Now divide the whole equation by dx whaddaya get? (no, am not latexing this for you ^_^)

OpenStudy (lgbasallote):

i have the integral thingies already...i need to know what those approaching means...for subbing

OpenStudy (anonymous):

\[ \large xy^2 dx + e^x dy = 0\\ \large xy^2 dx =-e^x dy \\ \large-\frac {dy}{y^2}= x e^{-x} dx\\ \large \frac 1 y = e^{-x}( 1-x) + C\\ \large y =\frac 1{ e^{-x}( 1-x) + C}\\ C=2,\, \text { from the condition given } \]

OpenStudy (anonymous):

I'm pretty sure, the right answer, is the right answer :s :3

OpenStudy (lgbasallote):

shouldnt there be a negative somewhere sir @eliassaab o.O

OpenStudy (anonymous):

The antidrevative of -1/y^2 is 1/y